# Thermal entanglement in conformal junctions

### v1

@@ Abstract
${}^{1}$ and INFN Sezione di Trieste, via Bonomea 265, 34136 Trieste, Italy.
${}^{2}$niversité Paris-Saclay, CNRS, LPTMS, 91405, Orsay, France.
${}^{3}$orbonne Université, CNRS, Laboratoire de Physique Théorique et Hautes Énergies, LPTHE, F-75005 Paris, France
@@ 1 Introduction
Single localized defects may alter macroscopically the behaviour of a quantum system. For one-dimensional critical quantum many-body systems, these effects are particularly enhanced, due to the presence of slow algebraic decay of correlations that propagate from the defect to the boundary. As a result, many counterintuitive phenomena have been observed in those systems, such as the famous Kondo effect [1].
@@
In the context of Luttinger liquids, it has been understood by Kane and Fisher [2], that, a single defect can be relevant, marginal or irrelevant. In the relevant case, which is particularly important for the Kondo Physics, the defect screens the correlation across it, and, in the infrared regime, the two halves of the systems become effectively decoupled. The underlying explanation for this phenomenon is nowadays formalized in the language of Boundary Conformal Field Theory (BCFT) [3, 4, 5].
@@
However, residual quantum correlations are still present at finite temperatures. Their quantitative understanding had been a challenge for a long time, but a recent work [6] shed light on their origin. In particular, the logarithmic negativity, introduced in Ref. [7] as a good measure of entanglement in mixed states, has been tackled in Ref. [6] via a non-trivial perturbation scheme around the infrared fixed point.
@@
Marginal defects are even more puzzling. In this case, the perturbation induced by the defect is invariant under the renormalization group, and novel infrared fixed points are found, corresponding to partially permeable interfaces. It has been shown that the ground state entanglement of a half chain attached to the defect grows logarithmically with the system size - with a universal prefactor which depends explicitly on the scattering properties of the defect [8, 9]. Since the prefactor above is related to the central charge for homogeneous system [10, 11], some literature has referred to an effective central charge in the presence of the interface. Besides this vague analogy, this terminology is highly misleading, since the bulk theory does not depend on the boundary condition, and, for instance, the associated central charge is not affected by the properties of the defect. Analytical investigation of the ground state entanglement for free systems (fermions and bosons) in the presence of interfaces has been pursued via Conformal Field Theory (CFT) techniques in Refs. [12, 13, 14], with lattice/microscopic methods in [15, 16, 17], and via the modular Hamiltonian in Ref. [18]. We also mention that linear growth of entanglement entropy after global quenches has been found in Refs. [19, 20].
@@
While extensive studies of thermal entanglement in CFT are present in the literature, see for example Refs. [21, 24, 25, 26, 27, 28, 29, 30, 31, 22, 23], and boundary effects have been thoroughly characterized [32, 33, 34, 35, 36, 37, 38, 39], few results are present for conformal defects. Importantly, a CFT approach to addressing this problem has been put forward in the work of Sakai and Satoh [9], in which the computation of the entropy is equivalently formulated as a partition function with the insertion of interface operators along the spatial direction. This approach, which has been used to characterize the ground state entropy between wires of the junction, turned out to be difficult to generalize to tackle other spatial bipartitions, thermal effects, excited states, dynamics, and so on.
@@
A solution to these issues has been proposed in [20], where the junction is formulated as a boundary CFT, and the entanglement is computed as a correlation function of certain twist fields. The novelty is that a family of twist fields, with a non-trivial boundary scaling dimension, has to be considered. For instance, these boundary scaling dimensions are precisely at the core of the unusual logarithmic growth of the ground state entropy, and they are related to the effective central charges observed in previous works.
@@
The aim of this work is to provide a qualitative and quantitative understanding of the thermal entanglement of conformal junctions (of $M$ wires), corresponding to marginal defects, via BCFT techniques, using the formalism of bulk/boundary twist fields. In particular, we characterize both the mutual information and the logarithmic negativity of spatial regions, as depicted in Fig. 1. To the best of our knowledge, both our approach and the results we found are new. For instance, the emerging physics strongly differs from the behaviors observed in Ref. [6] for relevant defects.
@@
In our analysis, we are able to get general closed analytical formulae describing the exact crossover as functions of the temperature. We also describe methods to tackle systematically the high/low-temperature regimes, using quantization over the direct/crossed channels respectively. We find that the entropies are, generically extensive with the subsystem size, as long as the typical geometrical lengths are much larger compared to the thermal length $\beta$, and the density of entropy does not depend on the details of the defect. This is physically reasonable since for the case considered above the dominant contribution to the entropy is due to classical correlations between the subsystem and a fictitious thermal bath. In contrast, both the mutual information and the logarithmic negativity of intervals attached to the defect satisfy an area law: the saturation value diverges logarithmically as the temperature decreases, and the universal prefactor is related to the boundary scaling dimensions of twist fields mentioned above. We finally mention that the same quantities remain finite for the relevant defect of Ref. [6], and they do not show any logarithmic divergence.
@@ 2 CFT approach: thermal effects on the half-line
In this section, we give a field theoretical description of a conformal junction of $M$ wires at finite temperature. We consider semi-infinite extended wires attached to a single defect, and we represent it as a theory over the half-plane with multiple species of fields and central charge $c$ a procedure known in the literature as the folded picture [12]. Since we work at finite temperature $\beta^{-1}$, the underlying Euclidean geometry is the half-infinite cylinder
@@
Periodic boundary conditions are chosen along the (Euclidean) time direction, while boundary conditions of type $b$, corresponding to the scattering properties of the defect, are kept along $\text{Re}\left(\zeta\right)=0$. We assume that $b$ is conformal-invariant, and information about the boundary condition is encoded in a conformal boundary state |$\ket{b}$ [40]. Such an objected is generically constructed using states from the Hilbert space of the corresponding bulk system, in our case a ring of length $\beta$ evolving in real time. We refer to this choice of quantization as the ”crossed” channel.
Our goal is to characterize the multi-point correlation functions of twist fields in the aforementioned geometry. These quantities play a crucial role in expressing various entanglement measures between spatial regions [11]. In principle, a theory contains numerous twist fields, each associated with different monodromy conditions with respect to the local fields [41, 42]. However, the forthcoming discussion is general and applies to any primary twist field, namely as long as it is the lightest and, therefore, the most relevant operator compatible with these specific semilocal properties.
Let us consider a twist field $\sigma(\zeta,\bar{\zeta})$inserted in the bulk and let $\tilde{\sigma}(\zeta,\bar{\zeta})$ be its hermitian conjugate. We assume that $\sigma(\zeta,\bar{\zeta})$ is a primary field with a given scaling dimension $\Delta$. In contrast, we refer to $\sigma_{b}$ as the corresponding boundary field arising from the leading term in the bulk-to-boundary OPE [43] expansion generated by the $\sigma$ operator  [^1] Compared to the $\beta=\infty$ case for which the relevant BCFT geometry is the half-plane, at finite temperature the boundary has non-zero intrinsic curvature. This requires the modification of the bulk-boundary OPE as explained in [44]. Conveniently, the leading term in (2.2) receives no curvature corrections, for $\sigma_{b}$ is primary. Also, the proportionality factors, known in the BCFT literature as bulk-boundary structure constants, have been discarded and absorbed in the normalization of the fields. In some contexts, they are nevertheless important, and, for example, they can be related to the Affleck-Ludwig boundary entropy [45]. However, they do not play a role in our discussion, as we will consider entanglement measures up to additive constants.
@@
where the (unit) proportionality constant has been absorbed in the normalization of the $\sigma_{b}$ field. Then, we employ the transformation law of primary fields [46] and we get
@@ 3 CFT approach: thermal effects at finite size
In this section, we explain how the approach of Sec. 2 can be adapted to deal with the finite size effects. This allows us to make a connection with the predictions of [9], where a conformal junction at finite size and zero temperature was considered. It also helps to understand the regimes where the presence of an additional boundary is relevant.
@@
which has the spectrum of a chiral CFT on a ring of length $2L$ (see Ref. [43] for additional details). We write the expectation value as
@@
From Eq. (3.4) it is evident that the first contribution is the zero temperature correlation function. Assuming a non-degenerate ground state, the first correction goes like $\sim e^{-\frac{\pi\beta}{L}\Delta_{\psi}}$, with $\Delta_{\psi}$ the dimension of the chiral field associated with the lightest excited state |$\ket{\psi}$. A direct calculation of ${}_{bb^{\prime}}\bra{0}\sigma_{b}(0)\tilde{\sigma}(\zeta,\bar{\zeta})\ket{0}_{bb^{\prime}}$, and $\bra{\psi}\sigma_{b}(0)\tilde{\sigma}(\zeta,\bar{\zeta})\ket{\psi}$ is nevertheless difficult. The reason is that the quantities above are, by conformal transformations, related to four-point functions in the upper half-plane, where $\sigma,\sigma_{b}$ and two boundary-changing operators are inserted. By Cardy’s doubling trick[40], this correlator is constructed with the conformal blocks of a $5$-point function on the complex plane, which are objects on which we do not have good analytic control.
@@ 3.1 Correlations between boundary twist fields
where a series in $e^{-\pi\frac{\beta}{L}}$ is present inside the parenthesis and $\Delta_{\psi}$ is the dimension of the lightest excited state. At strictly zero temperature and purely reflective boundary conditions $b^{\prime}$ at $\zeta=L$, one recovers the physical features discussed in Refs. [9, 12, 13, 14, 47], and the algebraic decay of the correlation function within $L$ corresponds to the ubiquitous logarithmic growth of entanglement entropy (see Sec. 4 for details). We also discover that setting permeable boundary conditions along $b^{\prime}$ leads to an additive contribution to entropy, at least at the leading order. Demonstrating this assertion at the microscopic level is challenging; for instance, the technique presented in Ref. [17] cannot be directly adapted. However, this remarkable outcome is a direct consequence of field theory, and a thorough investigation of its details is deferred to potential future studies.
@@
$\langle\sigma_{b}(0)\tilde{\sigma}_{b^{\prime}}(L)\rangle=\frac{{}_{\beta}\bra{b^{\prime}}\tilde{\sigma}_{b^{\prime}}(0)e^{-LH^{\prime}_{\text{twist.}}}\sigma_{b}(0)\ket{b}_{\beta}}{{}_{\beta}\bra{b^{\prime}}e^{-LH^{\prime}}\ket{b}_{\beta}}\simeq$ (3.13)
@@ 4 Entanglement measures
In this section, we discuss the relation between the thermal correlation functions of twist fields, characterized in Sec. 2, and the entanglement between spatial regions. We focus on the Rényi entropies [10] and logarithmic negativity [7] of a pair of intervals attached to a defect. We aim to understand the role of the boundary conditions at the defect, and how they affect the entanglement at finite temperature.
For simplicity, let us consider a conformal junction made of two semi-infinite wires connected together at a single point. This system is described by a (boundary) CFT denoted by [12]
@@
We parametrize the points of the junction via a pair[13]
@@
to parametrize the $n$ replicas. A family of twist fields in $\text{CFT}^{\otimes n}$ can be defined [^7] Strictly speaking, the twist fields are not local fields of the replica theory $\text{CFT}^{\otimes n}$. Still, an orbifold model $\text{CFT}^{\otimes n}/\mathbb{Z}_{n}$ can be constructed, promoting the replica permutation to global symmetry, and the twist fields would appear as local fields of the orbifold, as explained in detail in Ref. [48]. such that their insertion gives introduce branch cuts on each of the wires, implementing the appropriate replica shifts. For instance, following closely Ref. [20], we require that the twist fields [^8] We have opted to use a different notation for the twist fields of the junction orbifold CFT to stress that the discussion of this section and Sec. 5 is a particular case of the general results derived previously. $\mathcal{T}^{j},\tilde{\mathcal{T}}^{j}$ are the lightest fields satisfying
@@ 4.1 Rényi entropy
Here, we compute the Rényi entropies of $A,B$ and their union $A\cup B$. Given $\rho_{A}$ the reduced density matrix (RDM) of $A$, we express the $n$th Rényi entropy of $A$ as[10, 20]
@@
The bulk scaling dimension $\Delta$ depends only on the central charge $c_{1}$ of $\text{CFT}_{1}$, and its value is known [11]
@@
On the other hand, the boundary dimension $\Delta_{b}$, which characterizes the quantum fluctuation across the defect, depends non-trivially on the theory and the boundary conditions $b$: its explicit value is related to the logarithmic divergence of the ground state entropy with defect [20], that is known for free bosons and fermions (see e.g. Refs. [9, 49]). Since we want to keep the discussion general, we refrain from characterizing it further.
@@
which gives the mutual information [50] in the replica limit $n\rightarrow 1$.
@@
with $c_{1}+c_{2}$ the central charge of $\text{CFT}=\text{CFT}_{1}\otimes\text{CFT}_{2}$, while its boundary scaling dimension vanishes identically  [^9] In Ref. [11], and forthcoming literature, the insertion of the corresponding boundary twist field is even discarded, and it is formally expressed as the identity operator.. In particular, since $\Delta^{\prime}$ matches the sum of the dimensions of $\mathcal{T}^{1}$ and $\mathcal{T}^{2}$, no UV divergences arise by their fusion: this is not surprising, the two species of fields are decoupled in the bulk.
@@
As expected, the result of (4.15) is compatible with the prediction in [18] for the same bipartition, in the limit $\beta\gg\ell$ up to an overall constant which we did not evaluate in the present work.

### v2

# references.bib

@@ Abstract
1SISSA and INFN Sezione di Trieste, via Bonomea 265, 34136 Trieste, Italy.
2Université Paris-Saclay, CNRS, LPTMS, 91405, Orsay, France.
3Sorbonne Université, CNRS, Laboratoire de Physique Théorique et Hautes Énergies, LPTHE, F-75005 Paris, France
@@ 1 Introduction
Single localized defects may alter macroscopically the behaviour of a quantum system. For one-dimensional critical quantum many-body systems, these effects are particularly enhanced, due to the presence of slow algebraic decay of correlations that propagate from the defect to the boundary. As a result, many counterintuitive phenomena have been observed in those systems, such as the famous Kondo effect [Kondo-64].
@@
In the context of Luttinger liquids, it has been understood by Kane and Fisher [kf-92], that, a single defect can be relevant, marginal or irrelevant. In the relevant case, which is particularly important for the Kondo Physics, the defect screens the correlation across it, and, in the infrared regime, the two halves of the systems become effectively decoupled. The underlying explanation for this phenomenon is nowadays formalized in the language of Boundary Conformal Field Theory (BCFT) [al-91, al-93, oa-96].
@@
However, residual quantum correlations are still present at finite temperatures. Their quantitative understanding had been a challenge for a long time, but a recent work [kss-21] shed light on their origin. In particular, the logarithmic negativity, introduced in Ref. [Plenio-05] as a good measure of entanglement in mixed states, has been tackled in Ref. [kss-21] via a non-trivial perturbation scheme around the infrared fixed point.
@@
Marginal defects are even more puzzling. In this case, the perturbation induced by the defect is invariant under the renormalization group, and novel infrared fixed points are found, corresponding to partially permeable interfaces. It has been shown that the ground state entanglement of a half chain attached to the defect grows logarithmically with the system size - with a universal prefactor which depends explicitly on the scattering properties of the defect [Peschel-05, ss-08]. Since the prefactor above is related to the central charge for homogeneous system [cc-04, cc-09], some literature has referred to an effective central charge in the presence of the interface. Besides this vague analogy, this terminology is highly misleading, since the bulk theory does not depend on the boundary condition, and, for instance, the associated central charge is not affected by the properties of the defect. Analytical investigation of the ground state entanglement for free systems (fermions and bosons) in the presence of interfaces has been pursued via Conformal Field Theory (CFT) techniques in Refs. [gm-17, cmc-22, cmc-22a], with lattice/microscopic methods in [ep-10, isl-09, cmv-12], and via the modular Hamiltonian in Ref. [Mintchev:2020jhc]. We also mention that linear growth of entanglement entropy after global quenches has been found in Refs. [wwr-17, ce-23].
@@
While extensive studies of thermal entanglement in CFT are present in the literature, see for example Refs. [hn-13, dd-14, cw-14, ch-14, hn-15, cct-14, cw-15, hs-16, sr-19, wlckg-20, rmc-23], and boundary effects have been thoroughly characterized [bs-06, als-09, ntu-12, Takayanagi-11, bs-16, eir-22, eir-23, eirt-23], few results are present for conformal defects. Importantly, a CFT approach to addressing this problem has been put forward in the work of Sakai and Satoh [ss-08], in which the computation of the entropy is equivalently formulated as a partition function with the insertion of interface operators along the spatial direction. This approach, which has been used to characterize the ground state entropy between wires of the junction, turned out to be difficult to generalize to tackle other spatial bipartitions, thermal effects, excited states, dynamics, and so on.
@@
A solution to these issues has been proposed in [ce-23], where the junction is formulated as a boundary CFT, and the entanglement is computed as a correlation function of certain twist fields. The novelty is that a family of twist fields, with a non-trivial boundary scaling dimension, has to be considered. For instance, these boundary scaling dimensions are precisely at the core of the unusual logarithmic growth of the ground state entropy, and they are related to the effective central charges observed in previous works.
@@
The aim of this work is to provide a qualitative and quantitative understanding of the thermal entanglement of conformal junctions (of $M$ wires), corresponding to marginal defects, via BCFT techniques, using the formalism of bulk/boundary twist fields. In particular, we characterize both the mutual information and the logarithmic negativity of spatial regions, as depicted in Fig. 1. To the best of our knowledge, both our approach and the results we found are new. For instance, the emerging physics strongly differs from the behaviors observed in Ref. [kss-21] for relevant defects.
@@
In our analysis, we are able to get general closed analytical formulae describing the exact crossover as functions of the temperature. We also describe methods to tackle systematically the high/low-temperature regimes, using quantization over the direct/crossed channels respectively. Specifically, the sole required input consists of the boundary scaling dimension of the twist fields, which depends on the theory and the underlying defect. These dimensions correspond to the established prefactor of the logarithmic scaling of the ground state entropy, and we provide a relationship between their values and the entanglement measures by exploiting the conformal symmetry of the model. We find that the entropies are, generically extensive with the subsystem size, as long as the typical geometrical lengths are much larger compared to the thermal length $\beta$, and the density of entropy does not depend on the details of the defect. This is physically reasonable since for the case considered above the dominant contribution to the entropy is due to classical correlations between the subsystem and a fictitious thermal bath. In contrast, both the mutual information and the logarithmic negativity of intervals attached to the defect satisfy an area law: the saturation value diverges logarithmically as the temperature decreases, and the universal prefactor is related to the boundary scaling dimensions of twist fields mentioned above. We finally mention that the same quantities remain finite for the relevant defect of Ref. [kss-21], and they do not show any logarithmic divergence.
@@ 2 CFT approach: thermal effects on the half-line
In this section, we give a field theoretical description of a conformal junction of $M$ wires at finite temperature. We consider semi-infinite extended wires attached to a single defect, and we represent it as a theory over the half-plane with multiple species of fields and central charge $c$ a procedure known in the literature as the folded picture [gm-17]. Since we work at finite temperature $\beta^{-1}$, the underlying Euclidean geometry is the half-infinite cylinder
@@
Periodic boundary conditions are chosen along the (Euclidean) time direction, while boundary conditions of type $b$, corresponding to the scattering properties of the defect, are kept along $\text{Re}\left(\zeta\right)=0$. We assume that $b$ is conformal-invariant, and information about the boundary condition is encoded in a conformal boundary state |$\ket{b}$ [Cardy-84]. Such an objected is generically constructed using states from the Hilbert space of the corresponding bulk system, in our case a ring of length $\beta$ evolving in real time. We refer to this choice of quantization as the ”crossed” channel.
Our goal is to characterize the multi-point correlation functions of twist fields in the aforementioned geometry. These quantities play a crucial role in expressing various entanglement measures between spatial regions [cc-09]. In principle, a theory contains numerous twist fields, each associated with different monodromy conditions with respect to the local fields [bd-17, Dupic:2017hpb]. However, the forthcoming discussion is general and applies to any primary twist field, namely as long as it is the lightest and, therefore, the most relevant operator compatible with these specific semilocal properties.
Let us consider a twist field $\sigma(\zeta,\bar{\zeta})$inserted in the bulk and let $\tilde{\sigma}(\zeta,\bar{\zeta})$ be its hermitian conjugate. We assume that $\sigma(\zeta,\bar{\zeta})$ is a primary field with a given scaling dimension $\Delta$. In contrast, we refer to $\sigma_{b}$ as the corresponding boundary field arising from the leading term in the bulk-to-boundary OPE [Cardy-04] expansion generated by the $\sigma$ operator  [^1] Compared to the $\beta=\infty$ case for which the relevant BCFT geometry is the half-plane, at finite temperature the boundary has non-zero intrinsic curvature. This requires the modification of the bulk-boundary OPE as explained in [runkel2000boundary]. Conveniently, the leading term in (2.2) receives no curvature corrections, for $\sigma_{b}$ is primary. Also, the proportionality factors, known in the BCFT literature as bulk-boundary structure constants, have been discarded and absorbed in the normalization of the fields. In some contexts, they are nevertheless important, and, for example, they can be related to the Affleck-Ludwig boundary entropy [al-91a]. However, they do not play a role in our discussion, as we will consider entanglement measures up to additive constants.
@@
where the (unit) proportionality constant has been absorbed in the normalization of the $\sigma_{b}$ field. Then, we employ the transformation law of primary fields [dms-97] and we get
@@ 3 CFT approach: thermal effects at finite size
In this section, we explain how the approach of Sec. 2 can be adapted to deal with the finite size effects. This allows us to make a connection with the predictions of [ss-08], where a conformal junction at finite size and zero temperature was considered. It also helps to understand the regimes where the presence of an additional boundary is relevant.
@@
which has the spectrum of a chiral CFT on a ring of length $2L$ (see Ref. [Cardy-04] for additional details). We write the expectation value as
@@
From Eq. (3.4) it is evident that the first contribution is the zero temperature correlation function. Assuming a non-degenerate ground state, the first correction goes like $\sim e^{-\frac{\pi\beta}{L}\Delta_{\psi}}$, with $\Delta_{\psi}$ the dimension of the chiral field associated with the lightest excited state |$\ket{\psi}$. A direct calculation of ${}_{bb^{\prime}}\bra{0}\sigma_{b}(0)\tilde{\sigma}(\zeta,\bar{\zeta})\ket{0}_{bb^{\prime}}$, and $\bra{\psi}\sigma_{b}(0)\tilde{\sigma}(\zeta,\bar{\zeta})\ket{\psi}$ is nevertheless difficult. The reason is that the quantities above are, by conformal transformations, related to four-point functions in the upper half-plane, where $\sigma,\sigma_{b}$ and two boundary-changing operators are inserted. By Cardy’s doubling trick[Cardy-84], this correlator is constructed with the conformal blocks of a $5$-point function on the complex plane, which are objects on which we do not have good analytic control.
@@ 3.1 Correlations between boundary twist fields
where a series in $e^{-\pi\frac{\beta}{L}}$ is present inside the parenthesis and $\Delta_{\psi}$ is the dimension of the lightest excited state. At strictly zero temperature and purely reflective boundary conditions $b^{\prime}$ at $\zeta=L$, one recovers the physical features discussed in Refs. [ss-08, gm-17, cmc-22, cmc-22a, bcp-23], and the algebraic decay of the correlation function within $L$ corresponds to the ubiquitous logarithmic growth of entanglement entropy (see Sec. 4 for details). We also discover that setting permeable boundary conditions along $b^{\prime}$ leads to an additive contribution to entropy, at least at the leading order. Demonstrating this assertion at the microscopic level is challenging; for instance, the technique presented in Ref. [cmv-12] cannot be directly adapted. However, this remarkable outcome is a direct consequence of field theory, and a thorough investigation of its details is deferred to potential future studies.
@@
$\langle\sigma_{b}(0)\tilde{\sigma}_{b^{\prime}}(L)\rangle=\frac{{}_{\beta}\bra{b^{\prime}}\tilde{\sigma}_{b^{\prime}}(0)e^{-LH^{\prime}_{\text{twist.}}}\sigma_{b}(0)\ket{b}_{\beta}}{{}_{\beta}\bra{b^{\prime}}e^{-LH^{\prime}}\ket{b}_{\beta}}\simeq$
@@ 4 Entanglement measures
In this section, we discuss the relation between the thermal correlation functions of twist fields, characterized in Sec. 2, and the entanglement between spatial regions. We focus on the Rényi entropies [cc-04] and logarithmic negativity [Plenio-05] of a pair of intervals attached to a defect. We aim to understand the role of the boundary conditions at the defect, and how they affect the entanglement at finite temperature.
For simplicity, let us consider a conformal junction made of two semi-infinite wires connected together at a single point. This system is described by a (boundary) CFT denoted by [gm-17]
@@
We parametrize the points of the junction via a pair[cmc-22]
@@
to parametrize the $n$ replicas. A family of twist fields in $\text{CFT}^{\otimes n}$ can be defined [^7] Strictly speaking, the twist fields are not local fields of the replica theory $\text{CFT}^{\otimes n}$. Still, an orbifold model $\text{CFT}^{\otimes n}/\mathbb{Z}_{n}$ can be constructed, promoting the replica permutation to global symmetry, and the twist fields would appear as local fields of the orbifold, as explained in detail in Ref. [eir-22a]. such that their insertion gives introduce branch cuts on each of the wires, implementing the appropriate replica shifts. For instance, following closely Ref. [ce-23], we require that the twist fields [^8] We have opted to use a different notation for the twist fields of the junction orbifold CFT to stress that the discussion of this section and Sec. 5 is a particular case of the general results derived previously. $\mathcal{T}^{j},\tilde{\mathcal{T}}^{j}$ are the lightest fields satisfying
@@ 4.1 Rényi entropy
Here, we compute the Rényi entropies of $A,B$ and their union $A\cup B$. Given $\rho_{A}$ the reduced density matrix (RDM) of $A$, we express the $n$th Rényi entropy of $A$ as[cc-04, ce-23]
@@
The bulk scaling dimension $\Delta$ depends only on the central charge $c_{1}$ of $\text{CFT}_{1}$, and its value is known [cc-09]
@@
On the other hand, the boundary dimension $\Delta_{b}$, which characterizes the quantum fluctuation across the defect, depends non-trivially on the theory and the boundary conditions $b$: its explicit value is related to the logarithmic divergence of the ground state entropy with defect [ce-23], that is known for free bosons and fermions (see e.g. Refs. [ss-08, bb-15]). Since we want to keep the discussion general, we refrain from characterizing it further.
@@
which gives the mutual information [wvhc-08] in the replica limit $n\rightarrow 1$.
@@
with $c_{1}+c_{2}$ the central charge of $\text{CFT}=\text{CFT}_{1}\otimes\text{CFT}_{2}$, while its boundary scaling dimension vanishes identically  [^9] In Ref. [cc-09], and forthcoming literature, the insertion of the corresponding boundary twist field is even discarded, and it is formally expressed as the identity operator.. In particular, since $\Delta^{\prime}$ matches the sum of the dimensions of $\mathcal{T}^{1}$ and $\mathcal{T}^{2}$, no UV divergences arise by their fusion: this is not surprising, the two species of fields are decoupled in the bulk.
@@
As expected, the result of (4.15) is compatible with the prediction in [Mintchev:2020jhc] for the same bipartition, in the limit $\beta\gg\ell$ up to an overall constant which we did not evaluate in the present work.